OpenStudy (anonymous):

solve the compound inequality 5>-3x+4 or 7≤-4x+2

jhonyy9 (jhonyy9):

5-4>-3x so 3x>-1 so x> -1/3 7-2<=-4x so 4x<=-5 so x<= -5/4

OpenStudy (anonymous):

the optiond are (\[\infty\], -5/4 ] or (-1/3, \[\infty\]) or (-\[\infty\], \[\infty\]) or (-\[\infty\], 5/4 ] \[\upsilon\] (-1/3, \[\infty\])